Write the coefficient matrix of the system of equations below.
{5x-9y+5z=-8
{-6x-y-9z=-2
{2x+9y-7z=-2

Answer by stanbon(75887)

Write the coefficient matrix of the system of equations below.
{5x-9y+5z=-8
{-6x-y-9z=-2
{2x+9y-7z=-2
----
Just put the coefficients into a 3x3 matrix:
5...-9...5
-6..-1..-9
2....9..-7
====================
Cheers,
Stan H.